May 8, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and SeeVogh session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  • BCAL: Check max ratings for MPPCs (Elton).

Tentative Agenda

  1. Announcements
  2. Collaboration Meeting May 12-14
  3. Action Items
  4. FCAL
  5. BCAL
  6. Reconstruction and Simulation
  7. TDR update
  8. Calibration
  9. Any other business


Attendees: Manuel, Laura, Shayne, Will, Mark, Zisis, Elton, John L., David, Yi (JLab); Andrei, Tegan (UofR), Matt (IU).

  1. Announcements
    • CEBAF delivered 10.5 GeV beam to the tagger dump last night!
  2. Collaboration Meeting May 12-14
  3. Action Items
    • Elton: Note following meeting: Searched Hamamatsu literature for max voltage and could not find it in their documentation.
  4. FCAL
    • HDFCAL log book (John)
    • Fcal detector continues to work with no known problems
    • New firmware installed on Fadcs seems to have addressed all known problems with the Fadcs
    • Log book entry shows timing of LED pulser run, with the timing determined by the IU "linear interpolation" method
    • David: Note that the hardware this algorithm is not currently implemented in the hardware and plans for implementation are not clear
    • Timing of single events shows regions of similar timing and the pattern of timing is correlated with crates.
    • Problem is likely due to timing distribution to the crates
    • Pedestals: First look at overnight pedestal run shows some 3 hour structure (as for the BCAL) , but with much reduced amplitude (~0.6 channels)
  5. BCAL
    1. HDBCAL log book
    2. Pedestals (Mark)
      • First look at pedestal stability in BCAL
      • ~3 hour period for baseline oscillations with detector powered on
      • Same structure seen in data (not yet entered into the log) for pedestals with the power off, but at a reduced amplitude
      • Oscillations seem to correlate with the temperatures in the crates, which vary by about 2 degF. [Note: temperature variations of the MPPC chillers has a period of about 20 min and voltages are very stable.]
      • Matt: Each Fcal base has a temperature sensor that could be used as a reference for temperature variations in the hall. They accessible through the CAN controller that has a network address to access data.
      • Pedestals Changes in the pedestal when power is shut off to the MPPCs are about 20 counts. However, in one slot of one crate, they show variations of order 70-80 counts, which is likely not physical.
      • Studies are continuing to investigate
    3. BCAL Status (Will)
      • Orlando has written a script that cycles through 16 configurations of LED column and HV (4x4) for one quadrant at a time.
      • Will has developed an automatic procedure to check the response of each individual MPPC corresponding to a unique setting of LED and HV setting.
      • Plot in log book shows the results of the analysis of run 3030. Preliminary interpretation of the data indicate that there are some mis-wirings and half a dozen problems to be investigated.
    4. Status (Laura and Shayne)
      • Laura and Shayne have been checking the data by hand and see similar problems, but they are yet to do a direct comparison with Will
      • They also see indications of multiple LED columns firing when only one is expected.
  6. Reconstruction and Simulation
    • Last Thu and Tue Will, David, Zisis, Tegan and Mark met to establish the status of the BCAL reconstruction as a basis for Teagan to continue this work
    • Tegan (Zisis) have been in contact with Paul at CMU and Justin at MIT to develop the software further. He will go to MIT in July and work with Justin on BDTs
  7. TDR update
    • BCAL: Draft from Zisis has been updated by Elton. Mark read Elton's draft and provided useful comments. We will continue to include those into the next iteration that could be passed on to a larger group for comments.
    • FCAL: Matt is starting work on this section and will continue after the collaboration meeting.
  8. Calibration: no report
  9. Any other business